# Coding for Gaussian Two-Way Channels: Linear and Learning-Based Approaches

### v1

@@ III-F Optimization Solution for Sum-Error Minimization
We solved the weighted sum-power (WSP) minimization problem (in (29)) in Sec. III-D, and presented the solution algorithm, Algorithm 1. We can express the input-output relationship of Algorithm 1 as
@@
where $\eta_{1}$, $\eta_{2}$, and $\alpha$ are given. We note that the objective value in (29) is the sum of the transmit powers, where the transmit powers, $\mathbb{E}\big{[}\|{\bf x}_{1}\|^{2}\big{]}$ and $\mathbb{E}\big{[}\|{\bf x}_{2}\|^{2}\big{]}$, can be calculated from the obtained solutions, ${\bf g}_{1}$, ${\bf F}_{1}$, ${\bf g}_{2}$, and ${\bf F}_{2}$, by (16)-(17). To explicitly indicate the dependency of the powers, $\mathbb{E}\big{[}\|{\bf x}_{1}\|^{2}\big{]}$ and $\mathbb{E}\big{[}\|{\bf x}_{2}\|^{2}\big{]}$, on the values of $\eta_{1}$, $\eta_{2}$, and $\alpha$, we denote $\mathbb{E}\big{[}\|{\bf x}_{1}\|^{2}\big{]}$ and $\mathbb{E}\big{[}\|{\bf x}_{2}\|^{2}\big{]}$ as $P_{1}(\eta_{1},\eta_{2},\alpha)$ and $P_{2}(\eta_{1},\eta_{2},\alpha)$, respectively.
Our next goal is to solve the sum-error minimization problem in (25) by using the solution algorithm for the WSP minimization problem in (29), i.e., Algorithm 1. A naive approach is to first run Algorithm 1 for every feasible set of $(\eta_{1},\eta_{2},\alpha)$ and then find the best solution of $({\bf g}_{1},{\bf F}_{1},{\bf g}_{2},{\bf F}_{2})$ that minimizes the sum-error, $\mathcal{E}_{1}(\eta_{1})+\mathcal{E}_{2}(\eta_{2})$, and also satisfies the power constraints, $P_{i}(\eta_{1},\eta_{2},\alpha)\leq NP$, $i\in\{1,2\}$. However, the search space for $(\eta_{1},\eta_{2},\alpha)$ would be prohibitively large since the search space is continuous and three-dimensional.
@@ III-F1 Reducing the search space for $\alpha$ via the golden-section search method [22]
Given $\eta_{1}$ and $\eta_{2}$, we consider solving (29) for various $\alpha$ and obtaining the solution of (28) by selecting the best $\alpha$, denoted by $\alpha^{\star}$, and its corresponding solution, $({\bf g}_{1},{\bf F}_{1},{\bf g}_{2},{\bf F}_{2})$. However, searching over $\alpha$ requires lots of computations since $\alpha$ is continuous over the region $(0,1)$. Therefore, we develop an approximate searching method by assuming that the objective function in (28), $\max\big{\{}P_{1}(\eta_{1},\eta_{2},\alpha),P_{2}(\eta_{1},\eta_{2},\alpha)\big{\}}$, is unimodal along $\alpha\in(0,1)$. This assumption is motivated by our observation of unimodality through numerical experiments. Unimodal functions have a unique maximal or minimal point, for which the golden-section search method can be employed to find the point [22]. Thus, we adopt the golden-section search method and obtain
@@ III-F2 Reducing the search space for $\eta_{1}$ via the bisection method [25]
We recall that solving the original problem (25) is equivalent to solving (28) over all feasible region of ($\eta_{1}$, $\eta_{2}$) and finding the best solution. However, the computational complexity for searching over continuous two-dimensional space is prohibitive. Therefore, we use the fact that the objective function of (28) is increasing over $\eta_{1}$ from Lemma 1, and utilize a one-dimensional search method for finding $\eta_{1}$ given $\eta_{2}$. When seeking a particular functional value and its corresponding point in an increasing function, we can utilize the bisection method as an efficient one-dimensional search approach [25]. Hence, we apply the bisection method to find $\eta^{\star}_{1}(\eta_{2})$, which depends on the value of $\eta_{2}$, such that $u(\eta^{\star}_{1}(\eta_{2}))=0$, where
@@
For each value of $\eta_{2}$ in the search space, we obtain the tuple $(\eta_{1}^{\star}(\eta_{2}),\eta_{2},\alpha^{\star}(\eta_{1}^{\star},\eta_{2}))$ and find the best case that yields the minimum sum-error. Overall, by adopting the bisection method and golden-section search method in obtaining the solution of (25), we reduced the three-dimensional search space to one-dimensional search only over $\eta_{2}$ and mitigated the complexity. The overall algorithm is provided in Algorithm 2.
@@
9:Obtain the decoding schemes ${\bf w}_{1}$ and ${\bf w}_{2}$ by using ${\bf g}_{1}$, ${\bf F}_{1}$, ${\bf g}_{2}$, and ${\bf F}_{2}$ from (22).
@@ III-G Medium/Long Block-Lengths and Modulation Orders
We have focused on exchanging a message pair, $m_{1}$ and $m_{2}$, between the two users, where the encoding/decoding schemes are obtained by Algorithm 2. We recall that the message $m_{i}$ contains $K_{i}$ bits via $2^{K_{i}}$-ary PAM, $i\in\{1,2\}$. Let us define a long block-length as $L_{i}$ to distinguish from a short block-length $K_{i}$, where $K_{i}$ denotes the number of processing bits contained per each message $m_{i}$. Given a total of $L_{i}$ bits, User $i$ has a total of $M_{i}=L_{i}/K_{i}$ message symbols, each containing $K_{i}$ bits and modulated with $2^{K_{i}}$-ary PAM. We define the total number of channel uses for exchanging $L_{1}$ and $L_{2}$ bits between the two users as $N_{\text{L}}$ and the coding rate of User $i$ as $r_{i}=L_{i}/N_{\text{L}}$. We consider that the values of $L_{i}$ and $N_{\text{L}}$ are fixed, while the value of $K_{i}$ can be adjusted by User $i$. We assume the same number of messages at the users, i.e., $M=M_{1}=M_{2}$, by properly choosing $K_{i}$ such that $L_{1}/K_{1}=L_{2}/K_{2}$. The number of channel uses for exchanging a message pair is then $N=N_{\text{L}}/M$. Through this successive transmission, the users can exchange a total of $M$ message pairs over $N_{\text{L}}$ channel uses.
Depending on the value of $K_{i}$, there is a trade-off between the symbol distance and the feedback benefit in obtaining better SER in (24). If $K_{i}$ is small, e.g., $K_{i}=1$, the users only use a small number of channel uses for exchanging a message pair, which leads us not to fully benefit from the feedback scheme and thus to have a low SNR in (23). On the other hand, as $K_{i}$ becomes larger, the symbol error may easily occur since the distance between the quantized symbols decreases in signal space as the number of quantized symbols, $2^{K_{i}}$, increases. Therefore, a proper selection of $K_{i}$ or the modulation order $2^{K_{i}}$ is important to obtain better SER performance. We evaluate the error performance with different values of $K_{i}$ in Sec. VI.
@@
We have so far explored the linear approach for two-way coding, which offer low computational complexity (further elaborated in Sec. V). However, the simplified linear coding approach may limit the flexibility in code design. To address this, we propose a non-linear coding scheme based on deep learning in Sec. IV.
@@ IV-A State Propagation-Based Encoding
where we assume $y_{i}[0]=0$. Here, ${\bf s}_{i}[k]\in\mathbb{R}^{N_{s}}$ is the state vector, which propagates over time through the state-propagation function $h_{i}:\mathbb{R}^{K_{i}+N_{s}+1}\to\mathbb{R}^{N_{s}}$, which is given by
@@ IV-B1 GRUs for state propagation
We adopt two layers of unidirectional GRUs at each user to capture the time correlation of the receive signals in a causal manner. Formally, we represent the input-output relationship at each layer at time $k$ at User $i$ as
@@
Equations (59)-(60) can be represented as a functional form of the state propagation-based encoding in (56). By defining the overall state vector as ${\bf s}_{i}[k]=[{\bf s}_{i}^{(1)}[k],{\bf s}_{i}^{(2)}[k]]$, we obtain ${\bf s}_{i}[k]=h_{i}({\bf b}_{i},y_{i}[k-1],{\bf s}_{i}[k-1])$, where $h_{i}$ denotes the process of two layers of GRUs in (59)-(60). Note that ${\bf s}_{i}[k]$ propagates over time through the GRUs by incorporating the current input information into its state. Because the bit vector ${\bf b}_{i}$ with length $K_{i}$ is handled as a block to generate transmit signals with any length $N$, our method is flexible enough to support any coding rate $r_{i}=K_{i}/N$. Furthermore, the GRUs are interactive between the users since each user’s GRU incorporates the previously received symbol as input, as shown in Fig. 2. The pair of interactive GRUs effectively captures the interplay of the users’ encoding processes.
@@ IV-B3 Power control layer
The proof of Lemma 2 is provided in Appendix G.
@@ IV-D1 Model training
Either for sum-BLER or sum-BER minimization, we jointly train the encoders and decoders of both the users via autoencoder, where our overall coding architecture is provided in Fig. 4. The overall algorithm for training our RNN-based coding architecture is provided in Algorithm 3.
@@ IV-D2 Inference
(i) For sum-BLER minization, we define the loss function as
$\sum_{j\in\mathcal{I}_{\text{batch},t}}\mathcal{L}_{\text{sum-BLER},j}=-\sum_{j\in\mathcal{I}_{\text{batch},t}}\sum_{i=1}^{2}\sum_{\ell=1}^{2^{K_{i}}}{d}^{(j)}_{i}[\ell]\log\hat{d}^{(j)}_{i}[\ell],$ (70)
where ${d}_{i}^{(j)}[\ell]$ and $\hat{d}_{i}^{(j)}[\ell]$ are the $\ell$-th entry of ${\bf d}^{(j)}_{i}\in\{0,1\}^{2^{K_{i}}}$ and $\hat{\bf d}^{(j)}_{i}\in(0,1)^{2^{K_{i}}}$, respectively. Here, ${\bf d}_{i}^{(j)}=\text{one-hot}({\bf b}_{i}^{(j)})$ is the target vector, and $\hat{\bf d}_{i}^{(j)}$ is the block-level classification output obtained by the $j$-th data tuple passing through the overall encoder-decoder architecture in (59)-(67).
(ii) For sum-BER minization, we define the loss function as
$\sum_{j\in\mathcal{I}_{\text{batch},t}}\mathcal{L}_{\text{sum-BER},j}=-\sum_{j\in\mathcal{I}_{\text{batch},t}}\sum_{i=1}^{2}\sum_{\ell=1}^{K_{i}}\bigg{(}{b}_{i}^{(j)}[\ell]\log(\hat{d}^{(j)}_{i}[\ell])+(1-{b}_{i}^{(j)}[\ell])\log(1-\hat{d}^{(j)}_{i}[\ell])\bigg{)},$ (71)
where ${b}^{(j)}_{i}[\ell]$ and $\hat{d}_{i}^{(j)}[\ell]$ are the $\ell$-th entry of ${\bf b}^{(j)}_{i}\in\{0,1\}^{K_{i}}$ and $\hat{\bf d}^{(j)}_{i}\in(0,1)^{K_{i}}$, respectively. Here, ${\bf b}^{(j)}_{i}$ is the target vector, and $\hat{\bf d}_{i}^{(j)}$ is the bit-level classification output obtained by the $j$-th data tuple passing through (59)-(67).

### v2

@@ III-F Optimization Solution for Sum-Error Minimization
We solved the weighted sum-power (WSP) minimization problem  [^5] We solve it in a sub-optimal manner by trading off optimality for tractability. We note that optimality may be compromised due to the simplification of ${\bf F}_{2}$ in (47), the case when $\alpha<\frac{\sigma_{2}^{2}}{\sigma_{1}^{2}+\sigma_{2}^{2}}$ (shown in Corollary 1), and the iterative optimization process. The performance loss in WSP minimization is further analyzed in Sec. VI-J (in (46)) in Sec. III-D, and presented the solution algorithm, Algorithm 1. We can express the input-output relationship of Algorithm 1 as
@@
where $\eta_{1}$, $\eta_{2}$, and $\alpha$ are given. We note that the objective value in (46) is the sum of the transmit powers, where the transmit powers, $\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{1}\|^{2}\big{]}$ and $\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{2}\|^{2}\big{]}$, can be calculated from the obtained solutions, ${\bf g}_{1}$, ${\bf F}_{1}$, ${\bf g}_{2}$, and ${\bf F}_{2}$, by (16)-(17). To explicitly indicate the dependency of the powers, $\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{1}\|^{2}\big{]}$ and $\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{2}\|^{2}\big{]}$, on the values of $\eta_{1}$, $\eta_{2}$, and $\alpha$, we denote $\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{1}\|^{2}\big{]}$ and $\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{2}\|^{2}\big{]}$ as $P_{1}(\eta_{1},\eta_{2},\alpha)$ and $P_{2}(\eta_{1},\eta_{2},\alpha)$, respectively.
Our next goal is to solve the sum-error minimization problem in (45) (equivalently, (41)) by using the solution algorithm for the WSP minimization problem in (46), i.e., Algorithm 1. A naive approach is to first run Algorithm 1 for every possible combinations of $(\eta_{1},\eta_{2},\alpha)$, and then find the feasible solution set $\mathcal{S}$ in (45). Then, we select the best solution of $({\bf g}_{1},{\bf F}_{1},{\bf g}_{2},{\bf F}_{2})$ that minimizes the sum-error, $\mathcal{E}_{1}(\eta_{1})+\mathcal{E}_{2}(\eta_{2})$. However, the search space for $(\eta_{1},\eta_{2},\alpha)$ would be prohibitively large since the search space is continuous and three-dimensional.
@@ III-F1 Reducing the search space for $\alpha$ via the golden-section search method [25]
Given $\eta_{1}$ and $\eta_{2}$, we consider solving (46) for various $\alpha$ and obtaining the solution of (44) by selecting the best $\alpha$, denoted by $\alpha^{\star}$, and its corresponding solution, $({\bf g}_{1},{\bf F}_{1},{\bf g}_{2},{\bf F}_{2})$. However, searching over $\alpha$ requires lots of computations since $\alpha$ is continuous over the region $(0,1)$. Therefore, we develop an approximate searching method by assuming that the objective function in (44), $\max\big{\{}P_{1}(\eta_{1},\eta_{2},\alpha),P_{2}(\eta_{1},\eta_{2},\alpha)\big{\}}$, is unimodal along $\alpha\in(0,1)$. This assumption is motivated by our observation of unimodality through numerical experiments, which is demonstrated in the simulation section, Sec. VI-J. Unimodal functions have a unique maximal or minimal point, for which the golden-section search method can be employed to find the point [25]. Thus, we adopt the golden-section search method and obtain
@@ III-F2 Reducing the search space for $\eta_{1}$ via the bisection method [26]
We recall that the key in solving (45) is to find the feasible solution set $\mathcal{S}$, which requires solving (44) over all possible values of ($\eta_{1}$, $\eta_{2}$) and finding the best solution. However, the computational complexity for searching over continuous two-dimensional space is prohibitive. Therefore, we use the fact that the objective function of (44) is increasing over $\eta_{1}$ from Lemma 2, and utilize a one-dimensional search method for finding $\eta_{1}$ given $\eta_{2}$. Furthermore, we are interested in finding $\eta_{1}$ such that the max-power reaches the power constraint $NP$. If the max-power is less than $NP$, we can always find a higher value of $\eta_{1}$ by increasing the max-power to $NP$ given $\eta_{2}$ by Lemma 2. The higher value of $\eta_{1}$ (i.e., $\text{SNR}_{1}$) with the same value of $\eta_{2}$ will result in a lower sum-error value in (45). When seeking a particular functional value and its corresponding point in an increasing function, we can utilize the bisection method as an efficient one-dimensional search approach [26]. Hence, we apply the bisection method to find $\eta^{\star}_{1}(\eta_{2})$, which depends on the value of $\eta_{2}$, such that $u(\eta^{\star}_{1}(\eta_{2}))=0$, where
@@
For each value of $\eta_{2}$ in the search space, we obtain the tuple $(\eta_{1}^{\star}(\eta_{2}),\eta_{2},\alpha^{\star}(\eta_{1}^{\star},\eta_{2}))$ and find the best case that yields the minimum sum-error. Overall, by adopting the bisection method and golden-section search method in solving (45), we reduced the three-dimensional search space to one-dimensional search only over $\eta_{2}$ and mitigated the complexity. The overall algorithm is provided in Algorithm 2.
@@
9:Obtain the decoding schemes ${\bf w}_{1}$ and ${\bf w}_{2}$ by using ${\bf g}_{1}$, ${\bf F}_{1}$, ${\bf g}_{2}$, and ${\bf F}_{2}$ from (35) and (25).
@@ III-G Medium/Long Block-Lengths and Modulation Orders
We have focused on exchanging a message pair, $m_{1}$ and $m_{2}$, between the two users, where the encoding/decoding schemes are obtained by Algorithm 2. We recall that the message $m_{i}$ contains $K_{i}$ bits via $2^{K_{i}}$-ary PAM, $i\in\{1,2\}$. Let us define a long block-length as $L_{i}$ to distinguish from a short block-length $K_{i}$, where $K_{i}$ denotes the number of processing bits contained per each message $m_{i}$. Given a total of $L_{i}$ bits, User $i$ has a total of $M_{i}=L_{i}/K_{i}$ message symbols, each containing $K_{i}$ bits and modulated with $2^{K_{i}}$-ary PAM. We define the total number of channel uses for exchanging $L_{1}$ and $L_{2}$ bits between the two users as $N_{\text{L}}$ and the rate of User $i$ as $r_{i}=L_{i}/N_{\text{L}}$. We consider that the values of $L_{i}$ and $N_{\text{L}}$ are fixed, while the value of $K_{i}$ can be adjusted by User $i$. We assume the same number of messages at the users, i.e., $M=M_{1}=M_{2}$, by properly choosing $K_{i}$ such that $L_{1}/K_{1}=L_{2}/K_{2}$. The number of channel uses for exchanging a message pair is then $N=N_{\text{L}}/M$. Through this successive transmission, the users can exchange a total of $M$ message pairs over $N_{\text{L}}$ channel uses.
We provide new insights into how the value of $K_{i}$ affects sum-error rates in two-way channels, given that the rate $r_{i}=L_{i}/N_{\text{L}}$ is fixed, i.e., $L_{i}$ and $N_{\text{L}}$ are fixed. Specifically, depending on the value of $K_{i}$, there is a trade-off between the minimum distance of the constellation and the feedback benefit in obtaining better sum-error rates. First, as $K_{i}$ increases, the modulation order $2^{K_{i}}$ increases. Each messages is one of $2^{K_{i}}$ possible symbols, representing $K_{i}$ bits. For larger $K_{i}$, the minimum distance between the $2^{K_{i}}$ symbols decreases in signal space, making it more vulnerable to symbol errors. Second, as $K_{i}$ increases, the number of messages $M_{i}=L_{i}/K_{i}$ decreases, and the number of channel uses per message symbol $N=N_{\text{L}}/M_{i}=K_{i}/r_{i}$ increases. The increasing number of channel uses for exchanging a pair of messages results in greater benefit from feedback, leading to an improvement in error performance for the recipient of the help. This would result in better sum-error rates. In summary, as $K_{i}$ increases, the higher modulation order $2^{K_{i}}$ may lead to a loss in error performance by decreased minimum constellation distance, but the increased number of channel uses $N=K_{i}/r_{i}$ can improve sum-error performance due to greater feedback benefits. Therefore, sum-error performance varies depending on the values of $K_{i}$. The corresponding simulations are presented in Sec. VI-B.
@@
III-H Enhancement of the Linear Coding Scheme through Alternate Channel Uses
In two-way channels, it is important to utilize feedback at one user to minimize sum-errors. When the channel from User 2 to 1 is better than that of the reverse under certain conditions, User 2 acts as a helper to provide feedback to User 1 [6] and sends its message in the last channel use (by Proposition 3). In contrast, User 1 does not provide feedback to User 2 in this asymmetric channel (by Remark 3). Therefore, it can be understood that User 1 functions similarly to a transmitter as a recipient of help in feedback-enabled one-way channels [8, 10, 9, 15, 18], while User 2 acts as a receiver providing feedback to User 1. That is, to utilize feedback, User 1 waits for the transmit symbol to return, much like a transmitter in the feedback-enabled one-way channels. However, the number of channel uses for receiving feedback differs between feedback-enabled one-way and two-way channels. In feedback-enabled one-way channels, it takes only a single channel use for the transmitter to send the transmit symbols and receive feedback about the transmit symbols from the receiver. In contrast, in two-way channels, it takes two channel uses for each user to receive feedback about its transmit symbols. As a result, User 1 in two-way channels waits for the transmit symbol to return in two channel uses, which can lead to alternating channel uses.
We observe this pattern of alternate channel uses at the users in two-way channels from the solutions generated by Algorithm 2. Specifically, when exchanging a message pair, we observe that Users 1 and 2 only utilize about half of the channel uses (refer to the case of $N=3$ in Fig. 6(6(a)) and $N=5$ in Fig. 7(7(a))). More generally, when $N$ is odd, it has been shown that User 1 uses only odd-numbered channel uses, while User 2 uses only even-numbered channel uses and the last channel use. It is interesting to see that although we do not impose any constraints on the separation of the channel usages between the two users, the separation of the channel usages is caused by solving the optimization problem. Specifically, the solution ${\bf x}$ of the sub-problem (66) has non-zero and zero values in alternating positions.
It is important to note that the separation of the channel uses results in a waste of channel uses. To address this inefficiency, we introduce a strategy of using alternate channel uses for exchanging multiple pairs of messages, so that the two users fully utilize the available channel uses. For ease of explanation, we consider the example of exchanging $M=2$ message pairs, $(m_{1},m_{2})$ and $(m^{\prime}_{1},m^{\prime}_{2})$, over $N_{\text{L}}=6$ channel uses under power constraint $N_{\text{L}}P=6P$. The naive approach for exchanging the two message pairs is the successive transmission, as discussed in Sec. III-G; the first message pair is exchanged over the first $N=3$ channel uses under power constraint $3P$, and the second message pair is then exchanged over the next $N=3$ channel uses under power constraint $3P$. However, this approach results in a waste of channel uses. Therefore, we propose an alternate channel use strategy with the following steps.
${\bf x}_{1}=[a_{1},0,b_{1},0,c_{1}],\quad{\bf x}_{2}=[0,a_{2},0,b_{2},c_{2}],$ (73)
Note that the transmit signals in (75) satisfy the power constraint $N_{\text{L}}P=6P$ while utilizing $N_{\text{L}}=6$ channel uses. This strategy also satisfies causality of the encoding process. We note that the transmit signal $c_{2}^{\prime}$ in $\bar{\bf x}_{2}$ of (75) is not restricted to follow a causal processing since it is only composed of a scaled version of the message $m_{2}^{\prime}$ from Proposition 3, and thus it can be transmitted at any channel use. The further simulations are provided in Sec. VI-C
We have so far explored the linear approach for two-way coding, which offer low computational complexity (further elaborated in Sec. V). However, the simplified linear coding approach may limit the flexibility in code design. To address this, we propose a non-linear coding scheme based on deep learning in the next section.
@@ IV-A State Propagation-Based Encoding
where we assume $y_{i}[0]=0$. Here, ${\bf s}_{i}[k]\in\mathbb{R}^{N_{s}}$ is the state vector of size $N_{s}$, which propagates over time through the state-propagation function $h_{i}:\mathbb{R}^{K_{i}+N_{s}+1}\to\mathbb{R}^{N_{s}}$, which is given by
@@ IV-B1 GRUs for state propagation
We adopt two layers of uni-directional GRUs at each user to capture the time correlation of the receive signals in a causal manner. Formally, we represent the input-output relationship at each layer at time $k$ at User $i$ as
@@
Equations (80)-(81) can be represented as a functional form of the state propagation-based encoding in (77). By defining the overall state vector as ${\bf s}_{i}[k]=[{\bf s}_{i}^{(1)}[k],{\bf s}_{i}^{(2)}[k]]$, we obtain ${\bf s}_{i}[k]=h_{i}({\bf b}_{i},y_{i}[k-1],{\bf s}_{i}[k-1])$, where $h_{i}$ denotes the process of two layers of GRUs in (80)-(81). Note that ${\bf s}_{i}[k]$ propagates over time through the GRUs by incorporating the current input information into its state. Because the bit vector ${\bf b}_{i}$ with length $K_{i}$ is handled as a block to generate transmit signals with any length $N$, our method is flexible enough to support any rate $r_{i}=K_{i}/N$. Furthermore, the GRUs are interactive between the users since each user’s GRU incorporates the previously received symbol as input, as shown in Fig. 2. The pair of interactive GRUs effectively captures the interplay of the users’ encoding processes.
@@ IV-B3 Power control layer
The proof of Lemma 3 is provided in Appendix L.
@@ IV-D1 Model training
To minimize either the sum-BLER or sum-BER, we jointly train the encoders and decoders of both users within the structure of an autoencoder, as illustrated in Fig. 4. Specifically, the autoencoder framework allows us to optimize the encoder and decoder neural networks together within a single training process to achieve the common objective of minimizing the sum-BLER or sum-BER. The overall algorithm for training our RNN-based coding architecture is provided in Algorithm 3.
@@ IV-D2 Inference
The RNN-based coding architecture is trained to solve (78)-(79), which in turn produces distinct encoding/decoding modules at specific channel SNRs of the users. We consider the scenario where the users store multiple encoding/decoding architectures, each trained at distinct channel SNRs. During the inference stage, the users select and use a specific encoding/decoding architecture corresponding to the channel SNRs that the system configures. In this manner, the users adaptively replace their coding architectures depending on the measured channel SNRs.
@@
(i) For sum-BLER minization, we define the loss function as $\sum_{j\in\mathcal{I}_{\text{batch},t}}\mathcal{L}_{\text{sum-BLER},j}=-\sum_{j\in\mathcal{I}_{\text{batch},t}}\sum_{i=1}^{2}\sum_{\ell=1}^{2^{K_{i}}}{d}^{(j)}_{i}[\ell]\log\hat{d}^{(j)}_{i}[\ell],$ (91) where ${d}_{i}^{(j)}[\ell]$ and $\hat{d}_{i}^{(j)}[\ell]$ are the $\ell$-th entry of ${\bf d}^{(j)}_{i}\in\{0,1\}^{2^{K_{i}}}$ and $\hat{\bf d}^{(j)}_{i}\in(0,1)^{2^{K_{i}}}$, respectively. Here, ${\bf d}_{i}^{(j)}=\text{one-hot}({\bf b}_{i}^{(j)})$ is the target vector, and $\hat{\bf d}_{i}^{(j)}$ is the block-level classification output obtained by the $j$-th data tuple passing through the overall encoder-decoder architecture in (80)-(88).
(ii) For sum-BER minization, we define the loss function as $\sum_{j\in\mathcal{I}_{\text{batch},t}}\mathcal{L}_{\text{sum-BER},j}=-\sum_{j\in\mathcal{I}_{\text{batch},t}}\sum_{i=1}^{2}\sum_{\ell=1}^{K_{i}}\bigg{(}{b}_{i}^{(j)}[\ell]\log(\hat{d}^{(j)}_{i}[\ell])+(1-{b}_{i}^{(j)}[\ell])\log(1-\hat{d}^{(j)}_{i}[\ell])\bigg{)},$ (92) where ${b}^{(j)}_{i}[\ell]$ and $\hat{d}_{i}^{(j)}[\ell]$ are the $\ell$-th entry of ${\bf b}^{(j)}_{i}\in\{0,1\}^{K_{i}}$ and $\hat{\bf d}^{(j)}_{i}\in(0,1)^{K_{i}}$, respectively. Here, ${\bf b}^{(j)}_{i}$ is the target vector, and $\hat{\bf d}_{i}^{(j)}$ is the bit-level classification output obtained by the $j$-th data tuple passing through (80)-(88).